OpenStudy (blondey_315):

Find the direct variation equation of the graph through the points (0, 0) and (1, -2). Write in y=kx form. (Points : 1) y = 2x y = -2x y = 1/2 x y = -1/2 x

OpenStudy (surry99):

what have you tried so far?

OpenStudy (blondey_315):

nothing so far i'm still not sure how to do it could you help?

OpenStudy (surry99):

sure..how do you calculate the slope of a line if you have any two points on the line?

OpenStudy (surry99):

There is an equation for it...

OpenStudy (blondey_315):

hmm is it y2*x2 or somthing like that?

OpenStudy (surry99):

slope = (y2-y1)/(x2-x1)...can you now calculate the slope of the line?

OpenStudy (anonymous):

www.youtube.com/watch?v=JXRN_LkCa_o

OpenStudy (blondey_315):

hang on im doin it~

OpenStudy (blondey_315):

eh doubt i got i t right but i got 0...

OpenStudy (surry99):

ok...(0,0) is x1, y1 (1, -2) is x2, y2 (BTW you can reverse the selection of points, it makes no difference so slope = (y2 - y1)/(x2-x1) = (-2 -0)/(1-0) = -2 do you agree?

OpenStudy (blondey_315):

ooohhhhhh~ i see!

OpenStudy (surry99):

now from y = kx k is the slope you just calculated...

OpenStudy (surry99):

so the answer is.....?

OpenStudy (blondey_315):

so B? or y=-2x

OpenStudy (surry99):

yes...does that make sense to you know?

OpenStudy (blondey_315):

Yea i was looking at the equation wrong thank you for the help!

OpenStudy (surry99):

you are welcome...any time